# GALAX: Graph-Augmented Language Model for Explainable Reinforcement-Guided Subgraph Reasoning in Precision Medicine

### v1

@@ Abstract
In precision medicine, quantitative multi-omic features, topological context, and textual biological knowledge play vital roles in identifying disease-critical signaling pathways and targets, guiding the discovery of novel therapeutics and effective treatment strategies. Existing pipelines capture only one or two of these—numerical omics ignore topological context, text-centric LLMs lack quantitative grounded reasoning, and graph-only models underuse rich node semantics and the generalization power of LLMs—thereby limiting mechanistic interpretability. Although Process Reward Models (PRMs) aim to guide reasoning in LLMs, they remain limited by coarse step definitions, unreliable intermediate evaluation, and vulnerability to reward hacking with added computational cost. These gaps motivate jointly integrating quantitative multi-omic signals, topological structure with node annotations, and literature-scale text via LLMs, using subgraph reasoning as the principle bridge linking numeric evidence, topological knowledge and language context. To resolve this challenge, we propose GALAX (Graph Augmented LAnguage model with eXplainability), an innovative framework that integrates pretrained Graph Neural Networks (GNNs) into Large Language Models (LLMs) via reinforcement learning guided by a Graph Process Reward Model (GPRM), which generates disease-relevant subgraphs in a step-wise manner initiated by an LLM and iteratively evaluated by a pretrained GNN, enabling process-level supervision without explicit intermediate reasoning annotations. As an application, we also introduced Target-QA, a benchmark combining CRISPR-identified targets, multi-omic profiles, and biomedical graph knowledge across diverse cancer cell lines, which enables GNN pretraining for supervising step-wise graph construction and supports long-context reasoning over text-numeric graphs (TNGs), providing a scalable and biologically grounded framework for explainable, reinforcement-guided subgraph reasoning toward reliable and interpretable target and pathway discovery in precision medicine. The Target-QA [^1] Hugging Face: https://huggingface.co/datasets/FuhaiLiAiLab/Target-QA and GALAX [^2] GitHub: https://github.com/FuhaiLiAiLab/GALAX are publicly available at Huggingface and GitHub.
@@ 1 Introduction
Identifying therapeutic targets and elucidating disease mechanisms are main challenges in precision medicine (Steyaert et al., 2023; Topol, 2019). CRISPR-based gene editing has revolutionized functional genomic by enabling high-throughput perturbation of gene function across diverse cellular contexts (Shalem et al., 2014; Li et al., 2023). In oncology, large-scale CRISPR screens in cancer cell lines and patient-derived models have revealed context-specific genetic vulnerabilities, providing a robust experimental foundation for biomarker and target discovery (Shi et al., 2015). Despite these advances, computationally predicting key targets from multi-omic profiles and interpreting their mechanistic role in disease progression remains difficult. In particular, bridging omic data with interpretable explanations of molecular mechanisms continues to be a critical unmet need (Zhang et al., 2024a). Traditional approaches, such as differential expression analysis or essentiality scoring, lack the capacity to model the hierarchical and cross-modal dependencies in molecular networks, often overlooking key regulatory redundancies and pathway-level dynamics. Recent graph-based models have shown promise in outcome prediction tasks (Ren et al., 2024), yet they typically lack the structured supervision necessary for accurate target prioritization and mechanism discovery and seldom jointly integrate quantitative multi-omic features, topological structure with node annotations, and literature-scale text—limiting mechanistic interpretability.
@@
Meanwhile, Large Language Models (LLMs) have demonstrated strong capabilities in natural language understanding and reasoning, particularly through techniques like in-context learning (ICL) (Brown et al., 2020) and chain-of-thought (CoT) prompting (Wei et al., 2022), which enable multi-step reasoning. However, LLMs often suffer from hallucination and lack grounding in structured knowledge, especially in scientific domains. To mitigate these issues, Retrieval-Augmented Generation (RAG) (Lewis et al., 2020) and its graph-based variants, such as RoG (Luo et al., 2023) and G-Retriever (He et al., 2024), have been proposed to enhance LLM performance by incorporating external knowledge graph. Despite their utility, these methods primarily optimize for outcome-level accuracy, often neglecting the interpretability and reliability of intermediate reasoning steps. Moreover, they operate over purely textual graphs, limiting their ability to incorporate high-dimensional, personalized numerical data, such as multi-omic profiles, especially when the input context becomes large and complex, and and to respect biological topology at scale when the input context becomes large and complex. On the other hand, the Process Reward Model (PRM) framework has been introduced to provide fine-grained supervision over intermediate steps in reasoning tasks (Luo et al., ; Lightman et al., 2023; Uesato et al., 2022; Wang et al., 2023). PRMs provide step-wise supervision by assigning intermediate rewards to reinforcement learning (RL) agents, forming the foundation for Large Reasoning Models (LRMs) trained with Reinforcement Learning with Human Feedback (RLHF) (Bai et al., 2022), Proximal Policy Optimization (PPO) (Schulman et al., 2017), and Group Relative Policy Optimization (GRPO) (Shao et al., 2024). For example, StepGRPO (Zhang et al., 2025b) extends GRPO by incorporating rule-based step-wise rewards to supervise each intermediate reasoning step, addressing the sparse reward problem and enhancing multi-step reasoning in multimodal language models. However, PRMs face key limitations in defining fine-grained reasoning steps, verifying intermediate correctness, and hacking for model-based rewards (Gao et al., 2023), further complicating training.
@@
These challenges are amplified in biomedicine: reasoning over multi-omic, gene-regulatory text–numeric graphs (TNGs) lacks ground-truth stepwise annotations, making intermediate supervision infeasible, and the combinatorial explosion of biological paths renders exhaustive planning or retrieval impractical. We propose GALAX (Graph-Augmented LAnguage model with eXplainability), which couples LLMs with a pretrained GNN under reinforcement learning guided by a Graph Process Reward Model (GPRM). Instead of symbolic traces or manual labels, GALAX uses the GNN as a proxy stepwise supervisor, scoring intermediate subgraphs (partial signaling cascades) for biological plausibility and cancer relevance to provide fine-grained, graph-based rewards. GALAX prompts an LLM to propose candidate targets from multi-omic profiles and partial knowledge graphs, then an RL graph generator assembles task-specific cancer subnetworks under GPRM scoring—translating language reasoning into interpretable graph construction and yielding mechanistically grounded, patient-specific subnetworks for target prioritization. To evaluate, we introduce Target-QA, a benchmark integrating multi-omic data, biomedical graph knowledge, and CRISPR screening outcomes across diverse cancer cell lines. Together, GALAX and Target-QA deliver a scalable, reinforcement-guided solution for interpretable, patient-specific target identification and disease-mechanism discovery.
@@ 2 Related Work
LLMs Augmented with Knowledge and Graph Structures Prompt tuning has emerged as a lightweight and scalable method for adapting LLMs to downstream tasks without full finetuning (Korbak et al., 2023; Lester et al., 2021). While effective, it operates over flat text representations and struggles to incorporate structured domain knowledge or multi-modal signals. To address this, Retrieval-Augmented Generation (RAG) (Lewis et al., 2020), Chain-of-Thought (CoT) prompting (Wei et al., 2022), and graph-augmented approaches such as GraphRAG and G-Retriever (He et al., 2024) have been proposed to provide external knowledge and user-conditioned reasoning. RoG (Luo et al., 2023) further embeds relational graph information into prompts to improve performance. However, these methods rely unsupervised graph construction and do not support reliable reasoning. Furthermore, their reliance on purely textual inputs and limited context windows makes them ill-suited for integrating large-scale, patient-specific text-numeric graphs such as multi-omic signaling networks.
@@
Reinforcement Learning for Step-wise Reasoning Reinforcement learning has been instrumental in aligning LLM behavior through methods like RLHF (Bai et al., 2022), PPO (Schulman et al., 2017), and GRPO. The Process Reward Model (PRM) (Luo et al., ) enables step-wise supervision and has been adopted in Large Reasoning Models (LRMs). However, PRMs face key challenges: fine-grained step definitions are ambiguous, intermediate correctness is hard to validate, and model-based rewards can lead to reward hacking (Gao et al., 2023). These challenges are particularly acute in biomedical domains, where reasoning is inherently unstructured and lacks step-wise annotations, making conventional PRM pipelines impractical.
@@
Multi-omic Data Integration in Biomedical AI From a biological standpoint, the integration of genomic, transcriptomic, and proteomic has been essential for understanding disease mechanisms and therapeutic vulnerabilities (Hasin et al., 2017; Kristensen et al., 2014). Traditional approaches rely on statistical fusion or dimensionality reduction (Meng et al., 2016; Shen et al., 2009; Rohart et al., 2017; Argelaguet et al., 2018; Nguyen & Wang, 2020), which overlook the hierarchical and interconnected nature of molecular data. More recently, GNN-based models like MOGONET (Wang et al., 2021) and MoGCN (Li et al., 2022) have demonstrated the value of structured graph reasoning for cancer subtype classification and biomarker identification. However, these models are primarily designed for outcome prediction and often fall short in identifying actionable biomarkers associated with specific disease mechanisms.
@@
GALAX addresses the limitations of existing models by introducing a RL-guided framework that dynamically constructs biologically relevant subgraphs for each patient or cell line. This enables interpretable, context-sensitive target prioritization that adapts to both multi-omic features and disease-specific graph as the text-numeric format. To our knowledge, GALAX is the first to unify multi-omic integration, graph-augmented reasoning, and LLMs under a reinforcement learning paradigm with biologically grounded supervision, which learns to reason through step-wise subgraph generation, guided by an authoritative biomedical GRPM.
@@ 3 Problem Formulation
Identifying key targets and uncovering disease mechanisms remains a major challenge in precision medicine. To address this, we adopt the Text-Omic Signaling Graph (TOSG) (Zhang et al., 2025a), which integrates multi-omic features and biomedical context into a unified graph structure. Built upon TOSG, our model GALAX couples LLM-based hypothesis generation with reinforcement-guided subgraph reasoning to enable interpretable, patient-specific target predictions on the Target-QA benchmark.
@@
[…]
@@
We construct Target-QA using cancer cell lines with available CRISPR-based target information, each represented by multi-omic vector $X_{n}\in\mathbb{R}^{M}$, forming the set $\mathcal{X}=\{X_{1},X_{2},\cdots,X_{n},\cdots,X_{N}\}$. Due to input token constraints of LLMs (Achiam et al., 2023) and the presence of CpG sites with methylation beta values saturated at $1$, which dominate rankings and render top-$K$ selection uninformative, we selected the top $K$ features only from genomic, transcriptomic, and proteomic modalities to construct a concise multi-omic representation: $X_{n}^{(K)}=\{g_{1}^{\prime(n)},g_{2}^{\prime(n)},\dots,g_{K}^{\prime(n)},t_{1}^{\prime(n)},t_{2}^{\prime(n)},\dots,t_{K}^{\prime(n)},p_{1}^{\prime(n)},p_{2}^{\prime(n)},\dots,p_{K}^{\prime(n)}\}$, corresponding to omic-related graph nodes $\mathcal{V}_{n}^{(\text{omic})}$. Each cell line is additionally annotated with disease information curated from BioMedGraphica’s disease-target interaction graph $\mathcal{G}^{(\text{DTI})}=\{\mathcal{V}^{(\text{DTI})},\mathcal{E}^{(\text{DTI})}\}$, where $\mathcal{V}^{(\text{DTI})}=\{\mathcal{V}^{(S)},\mathcal{V}^{(p)}\}$ includes disease and protein nodes. And disease entity has has an associated name and textual description with $\mathcal{S}=\{S_{\text{name}},S_{\text{desc}}\}$ where $|\mathcal{S}|=M^{(S)}$. Cell lines and their disease labels form the sets $\mathcal{C}=\{c_{1},c_{2},\cdots,c_{n},\cdots,c_{N}\}$ and $\mathcal{S^{\prime}}=\{s^{\prime}_{1},s^{\prime}_{2},\cdots,s^{\prime}_{n},\cdots s^{\prime}_{N}\}$, respectively, based on annotations from DepMap (Dempster et al., ). To provide structured graph context, we apply a subgraph retrieval strategy that extracts disease-relevant protein entities $\mathcal{V}_{n}^{(p)}\subset\mathcal{V}^{(p)}$ and their interactions $\mathcal{E}^{(\text{DTI})}$, along with $h$-hop protein neighbors $\mathcal{V}_{n}^{(h)}$ by extracting interactions from $\mathcal{E}^{(\text{PPI})}$, where $\mathcal{V}_{n}^{(h)}\subset\mathcal{V}^{(p)}$. This yields a sample-specific subgraph $\mathcal{G}_{n}^{(\text{sub})}=\{\mathcal{V}_{n}^{(\text{sub})},\mathcal{E}_{n}^{(\text{sub})}\}$, where $\mathcal{V}_{n}^{(\text{sub})}=\mathcal{V}_{n}^{(p)}\cup{\mathcal{V}_{n}^{(h)}}$, with the full graph set $\mathcal{G}^{(\text{sub})}=\{\mathcal{G}_{1}^{(\text{sub})},\mathcal{G}_{2}^{(\text{sub})}\cdots,\mathcal{G}_{n}^{(\text{sub})},\cdots,\mathcal{G}_{N}^{(\text{sub})}\}$. Hence, each query $Q_{n}=\{c_{n},s^{\prime}_{n},X_{n}^{(K)},\mathcal{G}_{n}^{(\text{sub})}\}$ is paired with answer $A_{n}$ describing top-$\gamma$ CRISPR targets $R_{n}=\{r_{n,1},\dots,r_{n,\gamma}\}$, yielding instance $D_{n}=(Q_{n},A_{n})$ in dataset $\mathcal{D}=\{D_{1},\dots,D_{N}\}$, stratified by TCGA (Weinstein et al., 2013) types (e.g., $\mathcal{D}_{\text{LUAD}},\mathcal{D}_{\text{BRCA}}$).

### v2

@@ Abstract
In precision medicine, quantitative multi-omic features, topological context, and textual biological knowledge play vital roles in identifying disease-critical signaling pathways and targets, guiding the discovery of novel therapeutics and effective treatment strategies. Existing pipelines capture only one or two of these—numerical omics ignore topological context, text-centric LLMs lack quantitative grounded reasoning, and graph-only models underuse rich node semantics and the generalization power of LLMs—thereby limiting mechanistic interpretability. Although Process Reward Models (PRMs) aim to guide reasoning in LLMs, they remain limited by coarse step definitions, unreliable intermediate evaluation, and vulnerability to reward hacking with added computational cost. These gaps motivate jointly integrating quantitative multi-omic signals, topological structure with node annotations, and literature-scale text via LLMs, using subgraph reasoning as the principle bridge linking numeric evidence, topological knowledge and language context. To resolve this challenge, we propose GALAX (Graph Augmented LAnguage model with eXplainability), an innovative framework that integrates pretrained Graph Neural Networks (GNNs) into Large Language Models (LLMs) via reinforcement learning guided by a Graph Process Reward Model (GPRM), which generates disease-relevant subgraphs in a step-wise manner initiated by an LLM and iteratively evaluated by a pretrained GNN and schema-based rule check, enabling process-level supervision without explicit labels. As an application, we also introduced Target-QA, a benchmark combining CRISPR-identified targets, multi-omic profiles, and biomedical graph knowledge across diverse cancer cell lines, which enables GNN pretraining for supervising step-wise graph construction and supports long-context reasoning over text-numeric graphs (TNGs), providing a scalable and biologically grounded framework for explainable, reinforcement-guided subgraph reasoning toward reliable and interpretable target and pathway discovery in precision medicine. The Target-QA [^1] Hugging Face: https://huggingface.co/datasets/FuhaiLiAiLab/Target-QA and GALAX [^2] GitHub: https://github.com/FuhaiLiAiLab/GALAX are publicly available at Huggingface and GitHub.
@@ 1 Introduction
Identifying therapeutic targets and elucidating disease mechanisms are main challenges in precision medicine (Steyaert2023; Topol2019). CRISPR-based gene editing has revolutionized functional genomic by enabling high-throughput perturbation of gene function across diverse cellular contexts (shalem2014genome; Li2023-lx). In oncology, large-scale CRISPR screens in cancer cell lines and patient-derived models have revealed context-specific genetic vulnerabilities, providing a robust experimental foundation for biomarker and target discovery (shi2015discovery). Despite these advances, computationally predicting key targets from multi-omic profiles and interpreting their mechanistic role in disease progression remains difficult. In particular, bridging omic data with interpretable explanations of molecular mechanisms continues to be a critical unmet need (zhang2024using). Traditional approaches, such as differential expression analysis or essentiality scoring, lack the capacity to model the hierarchical and cross-modal dependencies in molecular networks, often overlooking key regulatory redundancies and pathway-level dynamics. Recent graph-based models have shown promise in outcome prediction tasks (ren2024survey), yet they typically lack the structured supervision necessary for accurate target prioritization and mechanism discovery and seldom jointly integrate quantitative multi-omic features, topological structure with node annotations, and literature-scale text—limiting mechanistic interpretability.
@@
Meanwhile, Large Language Models (LLMs) have demonstrated strong capabilities in natural language understanding and reasoning, particularly through techniques like in-context learning (ICL) (brown2020language) and chain-of-thought (CoT) prompting (wei2022chain), which enable multi-step reasoning. However, LLMs often suffer from hallucination and lack grounding in structured knowledge, especially in scientific domains. To mitigate these issues, Retrieval-Augmented Generation (RAG) (lewis2020retrieval) and its graph-based variants, such as RoG (luo2023reasoning), SubgraphRAG (li2024simple), GNN-RAG (mavromatis2025gnn) and G-Retriever (he2024g), have been proposed to enhance LLM performance by incorporating external knowledge graph. Despite their utility, these approaches still focus on final answer accuracy and give little attention to the reliability of intermediate reasoning. The retrieved subgraphs are noisy, large, and lack ground-truth mechanistic structure, making supervised retrieval unstable. Most existing models also fail to integrate numerical omic signals, causing the loss of cell line–specific information needed for target discovery. On the other hand, the Process Reward Model (PRM) framework has been introduced to provide fine-grained supervision over intermediate steps in reasoning tasks (luo2406improve; lightman2023let; uesato2022solving; wang2023plan). PRMs provide step-wise supervision by assigning intermediate rewards to reinforcement learning (RL) agents, forming the foundation for Large Reasoning Models (LRMs) trained with Reinforcement Learning with Human Feedback (RLHF) (bai2022training), Proximal Policy Optimization (PPO) (schulman2017proximal), and Group Relative Policy Optimization (GRPO) (shao2024deepseekmath). For example, StepGRPO (zhang2025r1) extends GRPO by incorporating rule-based step-wise rewards to supervise each intermediate reasoning step, addressing the sparse reward problem and enhancing multi-step reasoning in multimodal language models. However, PRMs face key limitations in defining fine-grained reasoning steps, verifying intermediate correctness, and hacking for model-based rewards (gao2023scaling), further complicating training.
@@
These challenges are amplified in biomedicine: reasoning over multi-omic, gene-regulatory text–numeric graphs (TNGs) lacks ground-truth stepwise annotations, making intermediate supervision infeasible, and the combinatorial explosion of biological paths renders exhaustive planning or retrieval impractical. We propose GALAX (Graph-Augmented LAnguage model with eXplainability), which couples LLMs with a pretrained GNN under reinforcement learning guided by a Graph Process Reward Model (GPRM). Instead of explicit labels, GALAX uses the GNN as a stepwise supervisor and schema-based rule term to check validity, scoring intermediate subgraphs (partial signaling cascades) for biological plausibility and cancer relevance to provide fine-grained, graph-based rewards. GALAX prompts an LLM to propose candidate targets from multi-omic profiles and partial knowledge graphs, then an RL graph generator assembles task-specific cancer subnetworks under GPRM scoring—translating language reasoning into interpretable graph construction and yielding mechanistically grounded, patient-specific subnetworks for target prioritization. To evaluate, we introduce Target-QA, a benchmark integrating multi-omic data, biomedical graph knowledge, and CRISPR screening outcomes across diverse cancer cell lines. Together, GALAX and Target-QA deliver a scalable, reinforcement-guided solution for interpretable, patient-specific target identification and disease-mechanism discovery.
@@ 2 Related Work
LLMs Augmented with Knowledge and Graph Structures Prompt tuning has emerged as a lightweight and scalable method for adapting LLMs to downstream tasks without full finetuning (korbak2023pretraining; lester2021power). While effective, it operates over flat text representations and struggles to incorporate structured domain knowledge or multi-modal signals. To address this, Retrieval-Augmented Generation (RAG) (lewis2020retrieval) and graph-augmented approaches such as G-Retriever (he2024g), RoG (luo2023reasoning), SubgraphRAG (li2024simple) and GNN-RAG (mavromatis2025gnn) have been proposed. However, these methods depend on accurate subgraph retrieval and still lack support for reliable reasoning. Most existing models also fail to integrate numerical omic signals, causing the loss of cell line–specific information needed for target discovery and leaving them poorly suited for large, patient-specific text–numeric graphs such as multi-omic signaling networks.
@@
Reinforcement Learning for Step-wise Reasoning Reinforcement learning has been instrumental in aligning LLM behavior through methods like RLHF (bai2022training), PPO (schulman2017proximal), and GRPO (shao2024deepseekmath). The Process Reward Model (PRM) (luo2406improve) enables step-wise supervision and has been adopted in Large Reasoning Models (LRMs). However, PRMs face key challenges: fine-grained step definitions are ambiguous, intermediate correctness is hard to validate, and model-based rewards can lead to reward hacking (gao2023scaling). These challenges are particularly acute in biomedical domains, where reasoning is inherently unstructured and lacks step-wise annotations, making conventional PRM pipelines impractical.
@@
Multi-omic Data Integration in Biomedical AI From a biological standpoint, the integration of genomic, transcriptomic, and proteomic has been essential for understanding disease mechanisms and therapeutic vulnerabilities (hasin2017multi; kristensen2014principles). Traditional approaches rely on statistical fusion or dimensionality reduction (meng2016dimension; shen2009integrative; rohart2017mixomics; argelaguet2018multi; nguyen2020multiview), which overlook the hierarchical and interconnected nature of molecular data. More recently, GNN-based models like MOGONET (wang2021mogonet) and MoGCN (li2022mogcn) have demonstrated the value of structured graph reasoning for cancer subtype classification and biomarker identification. However, these models are primarily designed for outcome prediction and often fall short in identifying actionable biomarkers associated with specific disease mechanisms.
@@
GALAX addresses the limitations of existing models by introducing a RL-guided framework that dynamically constructs biologically relevant subgraphs for each patient or cell line. This enables interpretable, context-sensitive target prioritization that adapts to both multi-omic features and disease-specific graph as the text-numeric format. To our knowledge, GALAX is the first to unify numerical multi-omic signals, literature-scale textual information, and biological topology under a reinforcement learning paradigm with biologically grounded supervision, which learns to reason through step-wise subgraph generation, guided by an authoritative biomedical GRPM.
@@ 3 Problem Formulation
Identifying key targets and uncovering disease mechanisms remains a major challenge in precision medicine. To address this, we adopt the Text-Omic Signaling Graph (TOSG) (zhang2025omnicelltosg), which integrates multi-omic features and biomedical context into a unified graph structure. Built upon TOSG, our model GALAX couples LLM-based hypothesis generation with reinforcement-guided subgraph reasoning to enable interpretable, patient-specific target predictions on the Target-QA benchmark.
@@
We construct the TOSG by integrating three modalities: numerical omic evidence $\mathcal{X}^{(0)}$, textual entity descriptions $\mathcal{T}$, and topological information within a unified biomedical knowledge graph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$. Here, $\mathcal{V}$ represents the topological entities and $\mathcal{E}$ denotes the biological relations.
@@
The detailed formulation of each modality is presented as follows. Feature Space: Specifically, utilizing the DepMap as the source of numerical evidence (dempster10extracting), we formulate this raw data as $\mathcal{X}^{(0)}=\{X_{n}^{(0)}\}_{n=1}^{N^{(0)}}$, where $X_{n}^{(0)}\in{\mathbb{R}^{{M}}}$ comprises of $N^{(0)}$ samples with $M$ entities. These samples are paired with binary labels for cancerous or non-cancerous cell lines, denoted as $\mathcal{Y}^{(0)}\in{\{0,1\}^{N^{(0)}}}$. Textual Space: Each entity in the TOSG is associated with a name and description, represented as $\mathcal{T}=\{T_{\text{name}},T_{\text{desc}}\}$, where $|T_{\text{name}}|=|T_{\text{desc}}|=M$. Topological Space: The $M$ entities in each sample are composed of multiple components derived from genomic, transcriptomic, and proteomic contexts. These multi-omic features are integrated into the TOSG using an existing integration tool, BioMedGraphica (zhang2024biomedgraphica), resulting in a text-attributed knowledge graph $\mathcal{G}=\{\mathcal{V,E}\}$. The set of vertices is defined as $\mathcal{V}=\{\mathcal{V}^{(pm)},\mathcal{V}^{(g)},\mathcal{V}^{(t)},\mathcal{V}^{(p)}\}$, representing promoter, gene, transcript, and protein entities, respectively. The size of each set is given by $|\mathcal{V}^{(pm)}|=m^{(pm)}$, $|\mathcal{V}^{(g)}|=m^{(g)}$, $|\mathcal{V}^{(t)}|=m^{(t)}$, and $|\mathcal{V}^{(p)}|=m^{(p)}$, such that $|\mathcal{V}|=m^{(pm)}+m^{(g)}+m^{(t)}+m^{(p)}=M$. Correspondingly, we map the cell-specific omic features for the $n$-th sample, denoted as $X^{(0)}_{n}\in\mathbb{R}^{M}$, using a compact vector concatenation of the four entity types: $X^{(0)}_{n}=\left[\mathbf{x}_{n}^{(pm)}\oplus\mathbf{x}_{n}^{(g)}\oplus\mathbf{x}_{n}^{(t)}\oplus\mathbf{x}_{n}^{(p)}\right]$, where $\mathbf{x}_{n}^{(\cdot)}$ represents the feature vector for the respective modality (e.g., $\mathbf{x}_{n}^{(p)}\in\mathbb{R}^{m^{(p)}}$ corresponds to protein levels). In detail, this graph can be decomposed into two subgraphs: $\mathcal{G}^{(\text{in})}=(\mathcal{V}^{(\text{in})},\mathcal{E}^{(\text{in})})$ and $\mathcal{G}^{(\text{PPI})}=(\mathcal{V}^{(\text{PPI})},\mathcal{E}^{(\text{PPI})})$. Here, $\mathcal{G}^{(\text{in})}$ captures the internal signaling processes for protein translation. As shown in Figure 2, internal propagation follows the central dogma (crick1970central): promoter (purple) $\to$ gene (red) $\to$ transcript (yellow) $\to$ protein (blue), with $\mathcal{V}=\mathcal{V}^{(\text{in})}$ and $|\mathcal{V}^{(\text{in})}|=M$. Meanwhile, $\mathcal{G}^{(\text{PPI})}$ represents the gene regulatory network structured around protein-protein interactions (PPI), where $\mathcal{V}^{(\text{PPI})}=\mathcal{V}^{(p)}$. In summary, the TOSG unifies numerical omic evidence, textual descriptions, and topological information into a Text-Numeric Graph, defined as $\mathcal{G}=\{\mathcal{X}^{(0)},\mathcal{T,V,E}\}$.
@@
We construct the feature set $\mathcal{X}$ by filtering for cancer cell lines that contain both comprehensive annotations and curated CRISPR-based target information, defining the collective multi-omic evidence as set $\mathcal{X}=\{X_{n}\}_{n=1}^{N}$, where $X_{n}\in{\mathbb{R}^{{M}}}$. Subsequently, to align this quantitative data with the reasoning capabilities of LLMs, we systematically structure the input context for each instance by integrating textual metadata, molecular profiles, and interaction graphs, detailed as follows. Omics Information: To augment context with patient-specific data, we incorporate the top-$K$ features for the LLM. Due to input token constraints of LLMs (achiam2023gpt) and the presence of CpG sites with methylation beta values saturated at $1$, which dominate rankings and render top-$K$ selection uninformative, we derive a concise multi-omic representation $X_{n}^{(K)}=[\mathbf{g}^{(K)}_{n}\oplus\mathbf{t}^{(K)}_{n}\oplus\mathbf{p}^{(K)}_{n}$] by extracting the top $K$ features from genomic, transcriptomic, and proteomic modalities. Here, $\mathbf{g}^{(K)}_{n},\mathbf{t}^{(K)}_{n}$, and $\mathbf{p}^{(K)}_{n}$ denote the ordered lists of selected gene, transcript, and protein names, respectively, which correspond directly to the omic-related graph nodes $\mathcal{V}_{n}^{(\text{omic})}$. Sample Information: Disease entity has an associated name and textual description with $\mathcal{S}=\{S_{\text{name}},S_{\text{desc}}\}$ where $|\mathcal{S}|=m^{(S)}$. Leveraging annotations from DepMap about cell lines, we mapped the cell line and disease names in samples to form the sets $\mathcal{C}=\{c_{n}\}_{n=1}^{N}$ and $\mathcal{S}^{\prime}=\{s^{\prime}_{n}\}_{n=1}^{N}$, respectively. Disease-related Protein Subgraph: For cell line, $c_{n}$, we use the cell line related disease entity $s^{\prime}_{n}$ to retrieve from BioMedGraphica’s disease-target interaction graph $\mathcal{G}^{(\text{DTI})}=\{\mathcal{V}^{(\text{DTI})},\mathcal{E}^{(\text{DTI})}\}$, where $\mathcal{V}^{(\text{DTI})}=\{\mathcal{V}^{(S)},\mathcal{V}^{(p)}\}$ includes disease and protein nodes. To provide structured graph context, we apply a subgraph retrieval strategy that extracts disease-relevant protein entities $\mathcal{V}_{n}^{(p)}\subset\mathcal{V}^{(p)}$ and their interactions $\mathcal{E}_{n}^{(\text{DTI})}$, along with $h$-hop protein neighbors $\mathcal{V}_{n}^{(h)}$ by extracting interactions from $\mathcal{E}^{(\text{PPI})}$, where $\mathcal{V}_{n}^{(h)}\subset\mathcal{V}^{(p)}$. This yields a sample-specific subgraph $\mathcal{G}_{n}^{(\text{sub})}=\{\mathcal{V}_{n}^{(\text{sub})},\mathcal{E}_{n}^{(\text{sub})}\}$, where $\mathcal{V}_{n}^{(\text{sub})}=\mathcal{V}_{n}^{(p)}\cup{\mathcal{V}_{n}^{(h)}}$, with the full graph set $\mathcal{G}^{(\text{sub})}=\{\mathcal{G}_{n}^{(\text{sub})}\}_{n=1}^{N}$. Hence, each query $Q_{n}=\{c_{n},s^{\prime}_{n},X_{n}^{(K)},\mathcal{G}_{n}^{(\text{sub})}\}$ is paired with answer $A_{n}$ describing top-$\gamma$ CRISPR targets $R_{n}=\{r_{n,1},\dots,r_{n,\gamma}\}$, yielding instance $D_{n}=(Q_{n},A_{n})$ in dataset $\mathcal{D}=\{D_{n}\}_{n=1}^{N}$, stratified by TCGA (weinstein2013cancer) types (e.g., $\mathcal{D}_{\text{LUAD}},\mathcal{D}_{\text{BRCA}}$).